# Automated Planning Domain Inference for Task and Motion Planning

### v2

@@ Abstract
Task and motion planning (TAMP) frameworks address long and complex planning problems by integrating high-level task planners with low-level motion planners. However, existing TAMP methods rely heavily on the manual design of planning domains that specify the preconditions and postconditions of all high-level actions. This paper proposes a method to automate planning domain inference from a handful of test-time trajectory demonstrations, reducing the reliance on human design. Our approach incorporates a deep learning-based estimator that predicts the appropriate components of a domain for a new task and a search algorithm that refines this prediction, reducing the size and ensuring the utility of the inferred domain. Our method is able to generate new domains from minimal demonstrations at test time, enabling robots to handle complex tasks more efficiently. We demonstrate that our approach outperforms behavior cloning baselines, which directly imitate planner behavior, in terms of planning performance and generalization across a variety of tasks. Additionally, our method reduces computational costs and data amount requirements at test time for inferring new planning domains.
@@ I Introduction
Robot autonomy that generalizes to diverse environments requires efficient integration of complex real-time perception, decision-making, and motion planning. Existing motion planning reaches its limits in complex and high-dimensional environments [Gonzalez2016-jl, Mohanan2018-ho, Paden2016-fa, Schwartz1988-ry, Zhou2022-vg] that are frequently encountered in modern robot applications [Liu2022-wj]. Task and motion planning (TAMP) frameworks address this challenge by employing high-level task planners to discretize a complex planning task into a sequence of manageable sub-tasks that low-level motion planners can complete.
While existing TAMP frameworks have solved many complex and long-horizon tasks in robotics [Garrett2021-ka, Lozano-Perez1979-zv], they rely heavily on human-designed planning domains, which are sets of logical rules and constraints that provide a basis for high-level task plans. Planning domains are restricted to predefined decision spaces and cannot easily be adapted to new tasks. This limitation drives the need for automated solutions that generate planning domains by leveraging information about existing, similar domains. We aim to develop a method that learns the relationships between tasks and their respective planning domains from training data, enabling the automatic generation of a new planning domain based on just one or a few demonstrations from humans.
This paper introduces an automated planning domain inference method that generates planning domains for new TAMP problems in a one-shot or few-shot manner, using demonstrations of continuous state-action trajectories. Our approach uses deep learning to predict domains for new tasks. To further improve the feasibility and planning efficiency of the planning domain, we combine these predictions with search algorithms to reduce the size of the planning domain to the smallest possible while ensuring that the generated domains are effective and reliable.
The primary contribution of this paper is (1) a graph attention network (GAT) trained on a small set of demonstrations to efficiently predict the planning domain for new tasks; (2) a guided generate-and-test search algorithm to refine the predicted planning domain, ensuring that planning is both feasible and efficient; (3) a framework that is deployable on a real robot that integrates GAT prediction and search to infer the planning domain for new tasks, requiring only a single human demonstration. We evaluate our method and compare it to existing baselines in 12 different environments. Our method significantly improves planning performance, generalizability, and computational costs.
Figure 1: Illustration of the general idea of this work: (1) One or a few human demonstrations for task execution are provided. 3D perception is used to extract object poses and interactions. (2) Plausible planning domains are inferred based on human demonstrations. (3) The generated domain is used for Task and Motion Planning and execution on real robots.
@@ II-A Learning to Plan for Faster Task and Motion Planning
Recent advances in machine learning have brought new approaches to task and motion planning. Learning-based TAMP methods often train policies that either accelerate or replace traditional planners. Recent research highlights the effectiveness of both supervised learning and reinforcement learning. Supervised learning methods demonstrate great ability to accelerate planning, but they often require extensive demonstration data and are conditioned to specific environments [Khodeir2023-so, Kim2020-du, Lin2022-tc, Silver2021-mv]. This limits their ability to generalize across different environment settings or goals [Dalal2023-cq, McDonald2021-uh, Zhu2021-sx, Yang2022-xa]. Reinforcement learning, as another major learning-based TAMP method in recent years, assists in accelerating the planning speed and enhances adaptability to dynamic and uncertain environments [Chitnis2016-pd, Jiang2019-gy, Paxton2017-lj, Xu2021-md]. However, creating a well-defined learning environment for complex TAMP problems, in simulation or in the real world, can be challenging. This reduces the applicability of reinforcement learning in TAMP scenarios [Wang2022-dx, Zhou2022-vg].
@@ II-B Learning Planning Domains
Recognizing the discussed limitations, researchers have increasingly shifted their focus toward learning planning domains. The mainstream learning-based methods in TAMP, designed to replicate the planner’s behavior, implicitly learns both the planning domain and the search algorithm. Focusing solely on learning the planning domain and integrating it with existing search algorithms reduces the burden on the learning algorithm.
Success in learning logical predicates from the environment  [Kase2020-ef, Mukherjee2021-bf, Silver2023-mi] facilitates the study of automatic generation of planning domains [Kumar2023-nk]. Automatic planning domain generation uses pre-existing logical predicates and actions to create planning domains autonomously. Recent studies have demonstrated that learning a planning domain enhances generalizability and adaptability compared to directly learning a TAMP planner [Diehl2021-qg]. However, existing methods encounter the challenges of high search costs and the requirement of a relatively large dataset [Kumar2023-nk, Silver2023-mi].
Our method is closely related to that of Kumar et al. [Kumar2023-nk], which identifies necessary preconditions and postconditions of an operator through a hill-climbing search. However, our method attempts to generate a full planning domain instead of merely operators (actions). Furthermore, rather than starting the search from scratch, our approach leverages transferable knowledge from existing planning domains, significantly accelerating the generation of domains and reducing the number of demonstrations required during execution. As a result, we are able to generate new planning domains for unseen tasks in a one-shot or few-shot manner.
Figure 2: The overall framework of domain inference. Left: The training phase of our method. Right: the testing phase of our method.
@@ III Problem setting
The standardized planning language used in this work is Planning Domain Definition Language (PDDL)  [McDermott1998PDDLthePD]. For a given task, the object set $\mathbf{O}=\langle o_{1},o_{2},\dots,o_{n}\rangle$ includes all relevant objects in the environment, with objects $o\in\mathbf{O}$ are references to entities in the planning problem [Khodeir2023-xs]. The state function $s:\mathbf{O}\times\mathbf{T}\rightarrow\mathcal{R}^{d\times n}$ provides the continuous state of an object, where $\mathbf{T}$ is the time indices denoting the time at which the state of an object is evaluated. This state is defined by continuous properties, such as pose and temperature.
@@
[…]
@@
Each logical action $\mathbf{\mathit{a}}\in\mathbf{A}$ has a corresponding precondition $\mathbf{Pre}=\langle\mathbf{\mathit{p_{1}}},\mathbf{\mathit{p_{2}}},...\rangle$, which is a set of predicates that must be true to trigger the action, and a postcondition $\mathbf{Eff}=\langle\mathbf{\mathit{p_{1}}},\mathbf{\mathit{p_{2}}},...\rangle$, which is a set of predicates that indicates the change in the logical environmental state after the action is executed. The transition between logical states before and after the actions is considered to be deterministic, $\mathbf{X_{\mathrm{before}}}\times\mathbf{\mathit{a}}\rightarrow\mathbf{X_{\mathrm{after}}}$. A task plan $\mathit{\pi}=\langle\mathbf{\mathit{a_{1}}},\mathbf{\mathit{a_{2}}},\mathbf{\mathit{a_{3}}},...,\mathbf{\mathit{a_{n}}}\rangle$ is a sequence of logical actions that change the initial logical environmental state to the goal logical environmental state. A (logical) trajectory $\mathbf{\tau}$ is the sequence of logical environmental state and logical action at each step during task execution, $\mathbf{\tau}=\langle(\mathbf{X_{1}},a_{1}),(\mathbf{X_{2}},a_{2}),\dots,(\mathbf{X_{n}},a_{n})\rangle$.
@@
For each $\mathbf{\mathit{a}}\in\mathbf{A}$, there is a corresponding $\mathbf{\mathit{c}}\in\mathbf{C}$ that verifies the feasibility of motion plans for the action and generates the plan if feasible. For example, the corresponding motion planner for the action ‘Pick’ will generate a collision-free trajectory for the robot to grasp an object and lift it from the table.
Using the concept defined above, we can now define a TAMP problem. Any TAMP problem $q$ can be defined by the object set, the full predicate set, the full action set, the initial logical environmental state, and the goal logical environmental state, $q=\langle\mathbf{O},\mathbf{P},\mathbf{A},\mathbf{X_{\mathrm{init}}},\mathbf{X_{\mathrm{goal}}},\rangle$. A TAMP problem set $\mathbf{Q}$ is a set of unsolved TAMP problems $\mathbf{Q}=\langle q_{1},q_{2},\dots\rangle$.
@@
A complete domain is a domain that includes all predicates and actions needed to solve the TAMP problem.
@@
An optimal domain is a complete domain that contains the least number of predicates and actions.
Given $\mathbf{A}$, $\mathbf{P}$, and $\mathbf{C}$, at the testing phase, our system aims at utilizing one (or a few) example trajectory $\mathbf{\tau_{e}}$ and a validation problem set containing $m$ unsolved TAMP problems, $\mathbf{Q}_{v}=\langle q_{1},q_{2},\dots,q_{m}\rangle$, to produce the optimal planning domain $\mathbf{D_{\mathrm{optm}}}=\langle\mathbf{P_{\mathrm{optm}}},\mathbf{A_{\mathrm{optm}}},\mathbf{C_{\mathrm{optm}}}\rangle$ that can reproduce $\mathbf{\tau_{e}}$ and generalize to similar TAMP tasks, even with shuffled object poses and varied object numbers.
@@ IV Methodology
More concretely, in Figure 2.I the training dataset includes information on TAMP problems with their associated planning domains. The predicate and action estimators output a relevance score $u\in[0,1]$ for each predicate and action, estimating their relevance of being part of the planning domain for a given problem $q$. The scores of predicates (actions) are denoted as $u_{\mathrm{p}}\in\mathit{U_{\mathrm{p}}}$ ($u_{\mathrm{a}}\in\mathit{U_{\mathrm{a}}}$), where the predicate (action) score set $\mathit{U_{\mathrm{p}}}$ ( $\mathit{U_{\mathrm{a}}}$) contains the score for all predicates (actions).
During testing, the inputs are an example trajectory $\mathbf{\tau_{e}}$ and a validation problem set $\mathbf{Q}_{v}$. As shown in Figure 2.II, the testing phase involves two main steps. First, the human demonstration (sensory data or human inputs) is converted into $\mathbf{\tau_{e}}$ to predict a distribution of relevant planning domains. Second, the predicted planning domains are refined by a search algorithm to find the optimal planning domain. If any of the domains explored during the search successfully solve all problems in the validation set $\mathbf{Q}_{v}$, the system returns the optimal domain $\mathbf{D_{\mathrm{optm}}}$. Otherwise, additional example trajectories of the same task are required.
@@ IV-A Predicate and Action Estimators
In Equation 1, $u_{\mathrm{p}}$ represents the score for a predicate, while $u_{\mathrm{a}}$ represents the score for an action. The scores for all possible domain sets can be computed using Equation 1. The domain set with the highest relevance score is denoted as $\omega_{\mathbf{D}}^{\mathrm{top}}$. The predicates and actions excluded from $\omega_{\mathbf{D}}^{\mathrm{top}}$ are ranked by their relevance score in a descending manner in a priority list $\mathbf{\mathit{L}}$. Similar score computation schemes can be found in learning-based incremental planning methodologies [Silver2021-mv].
@@ IV-B Dataset Generation
To effectively train the predicate and action estimator, we require a training dataset of TAMP problems and their associated planning domains. This dataset is generated using a sampling-based approach that randomly samples various planning problems in the simulation environment. Each TAMP problem is labeled with the relevant predicates and actions for their planning domain, assigning a score of 1 to those crucial for solving the problem and a score of 0 to irrelevant ones. The solvability of each TAMP problem is verified via a traditional TAMP planner  [Khodeir2023-so] deployed on a compute cluster.
Figure 3: Example on formulating logical states as scene graphs: (1) Unary predicates: Using the orange cube as an example, object ‘cube’ has an index of 1, (top, ?$o$), (cleaned, ?$o$) and (cooked, ?$o$) are all false. The feature node vector becomes $[1,0,0,0]$. We repeat this process for the other cubes to obtain the corresponding node features (2) Binary predicates: For example, the predicate (is_on, ?$o_{1}$, ?$o_{2}$ ) is true for the blue and orange cubes. Thus, an edge is formulated to connect them. Meanwhile, (is_on, ?$o_{3}$, ?$o_{2}$ ) is false, so there is no edge between green and orange cube.

### v3

@@ Abstract
Task and motion planning (TAMP) frameworks address long and complex planning problems by integrating high-level task planners with low-level motion planners. However, existing TAMP methods rely heavily on the manual design of planning domains that specify the preconditions and postconditions of all high-level actions. This paper proposes a method to automate planning domain inference from a handful of test-time trajectory demonstrations, reducing the reliance on human design. Our approach incorporates a deep learning-based estimator that predicts the appropriate components of a domain for a new task and a search algorithm that refines this prediction, reducing the size and ensuring the utility of the inferred domain. Our method can generate new domains from minimal test time demonstrations, enabling robots to handle complex tasks more efficiently. We demonstrate that our approach outperforms behaviour cloning baselines, which directly imitate planner behaviour, in terms of planning performance and generalization across a variety of tasks. Additionally, our method reduces computational costs and data amount requirements at test time for inferring new planning domains.
@@ I Introduction
Robot autonomy that generalizes to diverse environments requires efficient integration of complex real-time perception, decision-making, and motion planning. Existing motion planning algorithms reach their limits in complex and high-dimensional environments [10, 24, 26, 29, 36] that are frequently encountered in modern robot applications [19]. Task and motion planning (TAMP) frameworks address this challenge by employing high-level task planners to discretize a complex planning task into a sequence of manageable sub-tasks that low-level motion planners can complete.
While existing TAMP frameworks have solved many complex and long-horizon tasks in robotics [9, 20], they rely heavily on human-designed planning domains, which are sets of logical rules and constraints that provide a basis for high-level task plans. Planning domains are restricted to predefined decision spaces and cannot easily be adapted to new tasks. This limitation drives the need for automated solutions that generate planning domains by leveraging information about existing, similar domains. We aim to develop a method that learns the relationships between tasks and their respective planning domains from training data, enabling the automatic generation of a new planning domain based on just one or a few demonstrations from humans.
This paper introduces an automated planning domain inference method that generates planning domains for new TAMP problems in a one-shot or few-shot manner, using demonstrations of continuous state-action trajectories. Our approach uses deep learning to predict domains for new tasks. To further improve feasibility and planning efficiency, we combine these predictions with search to reduce the size of the planning domains, making them as small as possible while ensuring that they are effective and reliable.
The primary contribution of this paper is (1) a graph attention network (GAT) trained on a small set of demonstrations to efficiently predict the planning domain for new tasks; (2) a guided generate-and-test search algorithm to refine the predicted planning domain, ensuring that planning is both feasible and efficient; and (3) a framework that is deployable on a real robot that integrates GAT prediction and search to infer the planning domain for new tasks, requiring only a single human demonstration. We evaluate our method and compare it to existing baselines in 12 different environments. Our method significantly improves planning performance, generalizability, and computational costs.
Figure 1: Illustration of the general idea of our method: (1) One or a few human demonstrations for task execution are provided. 3D perception is used to extract object poses and interactions. (2) Plausible planning domains are inferred based on human demonstrations. (3) The generated domain is used for task and motion planning on a real robot.
@@ II-A Learning to Plan for Faster Task and Motion Planning
Recent advances in machine learning have brought new approaches to TAMP. Learning-based TAMP methods often train policies that either accelerate or replace traditional planners. Recent research highlights the effectiveness of both supervised learning and reinforcement learning. Supervised learning methods significantly accelerate planning, but they often require extensive demonstration data and are conditioned to specific environments [14, 15, 18, 30]. This limits their ability to generalize across different environment settings or goals [3, 23, 37, 35]. Reinforcement learning, as another major learning-based TAMP method in recent years, assists in accelerating planning speed and enhances adaptability to dynamic and uncertain environments [2, 11, 28, 34]. However, creating a well-defined learning environment for complex TAMP problems, in simulation or in the real world, can be challenging. This reduces the applicability of reinforcement learning in many TAMP scenarios [32, 36].
@@ II-B Learning Planning Domains
Recognizing the discussed limitations, researchers have increasingly shifted their focus toward learning planning domains. The mainstream learning-based methods in TAMP, designed to replicate the planner’s behaviour, implicitly learn both the planning domain and the search algorithm. Focusing solely on learning the planning domain and integrating it with existing search algorithms reduces the learning burden.
Success in learning logical predicates from the environment  [12, 25, 31] facilitates the study of automatic generation of planning domains [16]. Automatic planning domain generation relies on pre-existing logical predicates and actions to create planning domains autonomously. Recent studies have shown that learning a planning domain improves generalizability and adaptability compared to directly learning a TAMP planner [4]. However, existing methods encounter the challenges of high search costs and the requirement of a relatively large dataset [16, 31].
Our method is closely related to that of Kumar et al. [16], which identifies necessary preconditions and postconditions of an operator through a hill-climbing search. However, we attempt to generate a full planning domain instead of merely operators (actions). Furthermore, rather than starting the search from scratch, our approach leverages transferable knowledge from existing planning domains, significantly accelerating the generation of new domains and reducing the number of demonstrations required during execution. As a result, we are able to generate new planning domains for unseen tasks in a one-shot or few-shot manner.
Figure 2: Our overall framework for domain inference. Left: The training phase of our method. Right: the testing phase of our method.
@@ III Problem Setting
The standardized planning language used in our work is the Planning Domain Definition Language (PDDL) [22]. For a given task, the set of objects $\mathbf{O}=\langle o_{1},o_{2},\dots,o_{n}\rangle$ includes all relevant objects in the environment, where each object $o\in\mathbf{O}$ is a reference to an entity in the planning problem [13]. The state function $s:\mathbf{O}\times\mathbf{T}\rightarrow\mathcal{R}^{d\times n}$ provides the continuous state of an object, where $\mathbf{T}$ is the set of time indices at which the state of an object is evaluated. This state is defined by continuous properties, such as pose or temperature.
@@
Each predicate $\mathbf{\mathit{p}}\in\mathbf{P}$ defines a logical property, condition, or relationship among objects (e.g., (is_on $o_{1}$ $o_{2}$)) via a classifier function that outputs true or false given the continuous state $p(s(o_{1},t),s(o_{2},t),\dots)\in\{\text{true; false}\}$. A ground atom $x$ is a predicate combined with the objects and the classification result (e.g., (is_on $o_{1}$ $o_{2}$)). The logical state of an environment is a collection of all ground atoms for objects in the environment $\mathbf{X}=\langle\mathbf{\mathit{x_{1}}},\mathbf{\mathit{x_{2}}},\mathbf{\mathit{x_{3}}},\dots,\mathbf{\mathit{x_{n}}}\rangle$.
@@
Each logical action $\mathbf{\mathit{a}}\in\mathbf{A}$ has a corresponding precondition $\mathbf{Pre}=\langle\mathbf{\mathit{p_{1}}},\mathbf{\mathit{p_{2}}},$$\rangle$, which is a set of predicates that must be true to trigger the action, and a postcondition $\mathbf{Eff}=\langle\mathbf{\mathit{p_{1}}},\mathbf{\mathit{p_{2}}},$$\rangle$, which is a set of predicates that indicates the change in the logical environmental state after the action is executed. The transition between logical states before and after the actions is considered to be deterministic, $\mathbf{X_{\mathrm{before}}}\times\mathbf{\mathit{a}}\rightarrow\mathbf{X_{\mathrm{after}}}$. A task plan $\mathit{\pi}=\langle\mathbf{\mathit{a_{1}}},\mathbf{\mathit{a_{2}}},\mathbf{\mathit{a_{3}}},...,\mathbf{\mathit{a_{n}}}\rangle$ is a sequence of logical actions that change the initial logical environmental state to the goal logical environmental state. A (logical) trajectory $\mathbf{\tau}$ is the sequence of logical environmental state and logical action at each step during task execution, $\mathbf{\tau}=\langle(\mathbf{X_{1}},a_{1}),(\mathbf{X_{2}},a_{2}),\dots,(\mathbf{X_{n}},a_{n})\rangle$.
@@
For each $\mathbf{\mathit{a}}\!\in\!\mathbf{A}$, there is a corresponding $\mathbf{\mathit{c}}\!\in\!\mathbf{C}$ that verifies the feasibility of motion plans for the action and generates the plan if feasible. For example, the motion planner for the action ‘Pick’ will generate a collision-free trajectory for the robot to grasp an object and lift it from the table.
Using the concepts defined above, we can now define a TAMP problem. Any TAMP problem $q$ can be defined by the set of objects, the set of complete predicates, the set of full actions, the initial logical environmental state and the logical environmental state of the goal, $q=\langle\mathbf{O},\mathbf{P},\mathbf{A},\mathbf{X_{\mathrm{init}}},\mathbf{X_{\mathrm{goal}}},\rangle$. A TAMP problem set $\mathbf{Q}$ is a set of unsolved TAMP problems $\mathbf{Q}=\langle q_{1},q_{2},\dots\rangle$.
@@
A complete domain is a domain that includes all predicates and actions needed to solve a TAMP problem.
@@
An optimal domain is a complete domain that contains the least number of predicates and actions needed to solve a TAMP problem.
Given $\mathbf{A}$, $\mathbf{P}$, and $\mathbf{C}$, at the testing phase, our system aims at utilizing one (or a few) example trajectory $\mathbf{\tau_{e}}$ and a validation problem set containing $m$ unsolved TAMP problems, $\mathbf{Q}_{v}=\langle q_{1},q_{2},\dots,q_{m}\rangle$, to produce the optimal planning domain $\mathbf{D_{\mathrm{optm}}}=\langle\mathbf{P_{\mathrm{optm}}},\mathbf{A_{\mathrm{optm}}},\mathbf{C_{\mathrm{optm}}}\rangle$ that can reproduce $\mathbf{\tau_{e}}$ and generalize to similar TAMP tasks, even with shuffled object poses and varied object numbers. We assume $\mathbf{A}$ and $\mathbf{P}$ are given, as prior studies provide methods for generating predicates and actions [16, 31, 1, 17]. While these approaches do not learn complete planning domains from scratch, their findings form the basis of this work.
@@ IV Methodology
More concretely, in Figure 2 (I) the training dataset includes information on TAMP problems with their associated planning domains. The predicate and action estimators output a relevance score $u\in[0,1]$ for each predicate and action, estimating their relevance of being part of the planning domain for a given problem $q$. The scores of predicates (actions) are denoted as $u_{\mathrm{p}}\in\mathit{U_{\mathrm{p}}}$ ($u_{\mathrm{a}}\in\mathit{U_{\mathrm{a}}}$), where the predicate (action) score set $\mathit{U_{\mathrm{p}}}$ ( $\mathit{U_{\mathrm{a}}}$) contains the score for all predicates (actions).
During testing, the inputs are an example trajectory $\mathbf{\tau_{e}}$ and a validation problem set $\mathbf{Q}_{v}$. As shown in Figure 2 (II), the testing phase involves two main steps. First, the human demonstration (sensory data or human inputs) is converted into $\mathbf{\tau_{e}}$ to predict a distribution of relevant planning domains. Second, the predicted planning domains are refined by a search algorithm to find the optimal planning domain. If any of the domains explored during the search successfully solve all problems in the validation set $\mathbf{Q}_{v}$, the system returns the optimal domain $\mathbf{D_{\mathrm{optm}}}$. Otherwise, additional example trajectories of the same task are required.
@@ IV-A Predicate and Action Estimators
In Equation 1, $u_{\mathrm{p}}$ represents the score for a predicate, while $u_{\mathrm{a}}$ represents the score for an action. The scores for all possible domain sets can be computed using Equation 1. The domain set with the highest relevance score is denoted as $\omega_{\mathbf{D}}^{\mathrm{top}}$. The predicates and actions excluded from $\omega_{\mathbf{D}}^{\mathrm{top}}$ are ranked by their relevance score in a descending manner in a priority list $\mathbf{\mathit{L}}$. Similar score computation schemes can be found in learning-based incremental planning methodologies [30].
@@ IV-B Dataset Generation
To train the predicate and action estimator, we require a training dataset of TAMP problems and their associated planning domains. This dataset is generated using a sampling-based approach that randomly samples various planning problems in the simulation. Each TAMP problem is labeled with the relevant predicates and actions for their planning domain, assigning a score of 1 to those crucial for solving the problem and a score of 0 to irrelevant ones. The solvability of TAMP problems are verified via a traditional TAMP planner  [14] deployed on a compute cluster.
Figure 3: Example of formulating logical states as a scene graph: (1) Unary predicates: Using the orange cube as an example, object ‘cube’ has an index of 1, (top, ?o?o), (cleaned, ?o?o) and (cooked, ?o?o) are all false. The feature node vector becomes $[1,0,0,0]$. We repeat this process for the other cubes to obtain the corresponding node features (2) Binary predicates: For example, the predicate (is_on, $o_{1}$, $o_{2}$ ) is true for the blue and orange cubes. Thus, an edge is included to connect them. Meanwhile, (is_on, $o_{3}$, $o_{2}$ ) is false, so there is no edge between the green and the orange cube.